# An Efficient Continual Learning Framework for Multivariate Time Series Prediction Tasks with Application to Vehicle State Estimation

### v1

@@ IV Multivariate Representative Selection in time series tasks
After selecting the optimal representative set, it is stored in memory and used in conjunction with newly collected data during model training. This process helps to efficiently mitigate the effects of catastrophic forgetting in continual learning models. Algorithm 1 shows the step-by-step procedure of selecting representative set for memory-based continual learning in multi-output time series scenarios.
@@
Input: Historical datasets $D_{1},\,\dots,\,D_{k}$, and the newly collected dataset in new domain ($D_{k+1}$), where each $D_{i}$ represents the $i^{\text{th}}$ time series dataset, containing sensor measurements for inputs and corresponding target outputs.
@@
2Pre-train the model on historical dataset $D$.
@@
8            Within a fixed distance $d_{multi}$ for each captured point $y^{j}_{ig}$ in $i$, calculate the error between GLR values based on Eq. (23).
9            if $e<\text{th}$ then
@@
19Receive a dataset in new domain ($D_{k+1}$).
@@ VI-A Phase 1: Initial Training
In the initial step of the evaluation, we train the model parameters using the initial dataset. This step is common across all CL techniques, and the same trained model is used in the subsequent sections when new maneuvers are added. We focus on thoroughly training the model with the initial dataset to analyze the impact of forgetting on continual learning. As shown in Table I, the initial dataset comprises 10 maneuvers. The objective is to estimate both longitudinal and lateral velocities of the vehicle using input measurements and predicted outputs from previous steps (autoregressive structure) over the period of maneuver time.
@@ VI-A1 Feature Selection
To select the relevant inputs for estimating $V_{x}$ and $V_{y}$, in addition to the last predicted states of both $V_{x}$ and $V_{y}$ using the neural network model, we selected eight different signals: four wheel speeds, yaw rate, steering angle, and longitudinal and lateral accelerations, based on the vehicle model’s dynamic relation, as described in [25].
@@ VI-A2 Neural Network Model and Hyperparameter Set
This work adapts the original Transformer model to time series state estimation with the encoder-decoder mechanism. In the Encoder block, the multidimensional input data as selected in prevoius subsection, are fed into the model using a time-delay embedding (TDE) format based on a predefined window size. Let $p=8$ represent the dimensionality of each input time step and $d$ the time delay for the entire sequence. We considered a time delay of 0.5 seconds (50 measurements) for this experimental analysis. The input to the Decoder block consists of the last predicted output states of the model, allowing it to use previously predicted states alongside the output of the Encoder layer in the attention head. The decoder inputs are in a $q$-dimensional space, $q=2$ as we aim to estimate $V_{x}$ and $V_{y}$, with a time delay, which is set to match the time delay of the encoder inputs ($d$). To align with the input space, the decoder inputs are first embedded into the same dimension as the encoder inputs using a linear embedding layer, mapping the $q$-dimensional outputs to $p$ dimensions over all time sequences (time delays). Finally, the output of this block is processed through a feedforward layer with ReLU activation to introduce nonlinearities, followed by a linear layer that maps the preceding outputs to the final output layer, which has a dimension corresponding to the number of predicted outputs ($q=2$). The specifications of the hyperparameters for the Trasnformer model is outlined in Table II.
@@ VI-A3 Training and estimation results in initial training
During initial training, the model is stopped once the MSE loss falls below a predefined threshold. As previously mentioned, we use an autoregressive Transformer model to estimate each successive step of the maneuver. In this setup, the model feeds its prior output estimation into the input decoder rather than the actual output values, and estimation continues until the maneuver completes (typically 12 seconds in our experiments). For the test data, as noted, experiments were conducted within the same domain as the initial dataset to later assess the effect of catastrophic forgetting (CF) upon domain expansion. Table III presents the training time and estimation results on both the initial and test datasets, while Figure 5 illustrates the model’s estimation results on the test data within domain $D_{1}$ following initial training.
@@
As shown in the table and figure, the model has an acceptable performance in terms of estimation error, as the maneuver were performed in the same data domain as the initial dataset, there is no need to adapt the model to the test data.
@@ VI-B Phase 2: Continual Learning
Offline: Retrains all neural network parameters using both previously collected and newly acquired data, serving as a baseline or upper bound. This approach, referred to as Offline or Joint training mode, follows the terminology used in [26].
@@
We consider a specified number of points as memory points for A-GEM ($m_{1}=150$), using the same quantity for EM-ReSeleCT. However, the key difference in EM-ReSeleCT is the selection of a specific representative set rather than random selection. This multivariate analysis within EM-ReSeleCT’s structure significantly reduces the number of memory points required. In our approach, after processing historical data with EM-ReSeleCT, the total number of memory points is minimized to $m_{1}=97$ memory points.
Table IV presents the training time and Mean Absolute Error (MAE) of each continual learning strategy on the new dataset after introducing the domain into the model. As shown, all strategies demonstrate strong performance in estimation error for the new maneuver. In terms of training time, the Offline mode requires substantial time due to its use of all historical data combined with the new data during training. Interestingly, both A-GEM and EM-ReSeleCT have comparable training times, despite EM-ReSeleCT’s higher computational demands. This efficiency is attributed to the significantly reduced final memory set in EM-ReSeleCT compared to A-GEM or random selection strategies. SI and None modes have the same training time, as neither retains memory points in the continual learning phase.
@@
A critical aspect of continual learning techniques is their capacity to preserve previously learned knowledge across domains encountered so far. Table V and Figures 7 and 8 display the estimation results on the test data (in $D_{1}$) using different CL strategies, following the introduction of a new domain (in $D_{2}$) to the model. As shown, the proposed method (EM-ReSeleCT) achieves an estimation error comparable to that of the Offline mode, which serves as our baseline. EM-ReSeleCT significantly outperforms other popular CL strategies, such as A-GEM and SI. This superior performance over A-GEM is attributed to EM-ReSeleCT’s capability to select informative memory points rather than relying on random selection, coupled with optimization modifications that enable positive backward transfer, enhancing learning on previously encountered data. SI, on the other hand, shows poor performance, as weight regularization methods have consistently demonstrated limited accuracy in these regression problems.
@@
Figure 7 presents the estimation results of EM-ReSeleCT in comparison to the Offline and None modes. As illustrated, for both longitudinal ($V_{x}$) and lateral ($V_{y}$) velocities, the proposed method demonstrates promising performance, closely aligning with the Offline mode but with significantly lower computational time. In specific regions, EM-ReSeleCT effectively preserves information, whereas in the None mode—where prior information is disregarded—the model performs poorly.
@@
As previously mentioned, EM-ReSeleCT enhances A-GEM’s optimization algorithm to better preserve historical data, as outlined in Eq. (10). Figure 9 illustrates the loss on memory points across training epochs in a continual learning scenario. Our algorithm not only prevents an increase in memory point loss but also reduces it over certain epochs, demonstrating improved retention of historical information. In addition to the modified optimization, EM-ReSeleCT selects key information by identifying informative data points rather than using A-GEM’s random selection, which further mitigates catastrophic forgetting across all historical data. Figure … shows the loss error on the original historical data over each training epoch. As shown, EM-ReSeleCT significantly outperforms A-GEM, as it preserves the most vulnerable memory points—those most susceptible to catastrophic forgetting.
@@ VI-D Uncertainty Quantification with Conformal Prediction
$q_{1-\alpha}=\text{Quantile}_{1-\alpha}\{r_{1},r_{2},\dots,r_{n}\}.$
@@
$\text{PI}(x_{\text{test}})=\left[f(x_{\text{test}})-q_{1-\alpha},f(x_{\text{test}})+q_{1-\alpha}\right].$
This interval is guaranteed to cover the true target value $y_{\text{test}}$ with probability at least $1-\alpha$, under the assumption that the calibration set is exchangeable with the test data. In this study, we applied conformal prediction to estimate the prediction intervals for the outputs $V_{x}^{\text{SM}}$ and $V_{y}^{\text{SM}}$. Separate quantiles were computed for each output dimension, providing tailored uncertainty bounds for multivariate predictions.
To ensure reliable uncertainty quantification, we recalibrate the CP model after introducing the new dataset. The calibration data includes samples from both the old and new datasets to ensure comprehensive coverage. After determining the quantile $q_{\alpha}$ for each model, we apply the uncertainty quantification to both the test data from the old data space and the new data space.
Figure 10 a and b, show the coverage rate and interval width for the estimation of $V_{x}$ with different numbers of memory points selected for the proposed model on the test dataset from the old data space.
Figure 10 c and d, illustrate the coverage rate and interval width for the estimation of $V_{y}$ with different numbers of memory points selected for the proposed model.
As illustrated in the figures, the interval size and uncertainty level increase significantly as the number of memory points decreases. Furthermore, the figures demonstrate that after increasing the number of memory points beyond a certain threshold, the interval size remains unchanged. This suggests that storing additional memory points beyond this threshold does not contribute to further efficiency gains.
Figure 10: Uncertainty analysis of different models by number of selected memory points, on the old (test) dataset. a and b: covcerage rate and interval width vs the number of memory points on $Vx$, respectively. c and d: covcerage rate and interval width vs the number of memory points on $Vy$, respectively.
@@ VII Conclusions and Future Work
In this paper, we introduce EM-ReSeleCT, an efficient multivariate representative selection and optimization strategy for continual learning in time series tasks. Experimental analysis on multivariate output estimation using Transformer models, applied to the electric Equinox vehicle, demonstrates the superior performance of EM-ReSeleCT in both training time and estimation accuracy compared to other state-of-the-art continual learning algorithms. Specifically, EM-ReSeleCT achieves estimation errors close to the Offline mode while significantly reducing computational time. The method effectively minimizes the number of memory points by identifying instances where two selected points across outputs represent the same information but are in close proximity.

### v2

@@ IV Multivariate Representative Selection in time series tasks
After selecting the optimal representative set, it is stored in memory and used in conjunction with newly collected data during model training. This process efficiently mitigates the effects of catastrophic forgetting in continual learning models. Algorithm 1 shows the step-by-step procedure for selecting a representative set for memory-based continual learning in multi-output time series scenarios.
@@
Input: Historical datasets $D_{1},\,\dots,\,D_{k}$, and the newly collected dataset in new domain $D_{k+1}$, where each $D_{i}$ represents the $i^{\text{th}}$ time series dataset, containing sensor measurements for inputs and corresponding target outputs.
@@
2Pre-train the model on the historical dataset $D$.
@@
8            Within a fixed distance $d_{multi}$ for each captured point $y^{j}_{ig}$ in $i$, calculate the GLR difference based on Eq. (18).
9            if $E<\text{threshhold}$ then
@@
19Receive a dataset in the new domain $D_{k+1}$.
@@ VI-A Phase 1: Initial Training
In the initial step of the evaluation, we train the model parameters using the initial dataset. This step is common across all continual learning techniques, and the same trained model is used in the subsequent sections when new maneuvers are added. We focus on thoroughly training the model with the initial dataset to analyze the impact of forgetting on continual learning. As shown in Table I, the initial dataset comprises 10 maneuvers. The objective is to estimate both longitudinal and lateral velocities of the vehicle using input measurements and predicted outputs from previous steps (autoregressive structure) over the period of maneuver time.
@@ VI-A1 Feature Selection
To select the relevant inputs for estimating longitudinal ($V_{x}$) and lateral ($V_{y}$) velocities, in addition to the last predicted states of both $V_{x}$ and $V_{y}$ using the neural network model, we selected eight different signals: four wheel speeds, yaw rate, steering angle, and longitudinal and lateral accelerations, based on the vehicle model’s dynamic relation, as described in [25].
@@ VI-A2 Neural Network Model and Hyperparameter Set
This work adapts the original Transformer model to time series state estimation with the encoder-decoder mechanism. In the Encoder block, the multidimensional input data as selected in the previous subsection is fed into the model using a time-delay embedding (TDE) format based on a predefined window size. Let $p=8$ represent the dimensionality of each input time step and $d$ the time delay for the entire sequence. We considered a time delay of 0.5 seconds (50 measurements) for this experimental analysis. The input to the Decoder block consists of the last predicted output states of the model, allowing it to use previously predicted states alongside the output of the Encoder layer in the attention head. The decoder inputs are in a $q$-dimensional space, $q=2$ as we aim to estimate $V_{x}$ and $V_{y}$, with a time delay, which is set to match the time delay, $d$, of the encoder inputs. To align with the input space, the decoder inputs are first embedded into the same dimension as the encoder inputs using a linear embedding layer, mapping the $q$-dimensional outputs to $p$ dimensions over all time sequences (time delays). Finally, the output of this block is processed through a feedforward layer with ReLU activation to introduce nonlinearities, followed by a linear layer that maps the preceding outputs to the final output layer, which has a dimension corresponding to the number of predicted outputs ($q=2$). The specifications of the hyperparameters for the Transformer model are outlined in Table II.
@@ VI-A3 Training and estimation results in initial training
During initial training, the model is stopped once the MSE loss falls below a predefined threshold. As previously mentioned, we use an autoregressive Transformer model to estimate each successive step of the maneuver. In this setup, the model feeds its prior output estimation into the input decoder rather than the actual output values, and estimation continues until the maneuver completes (typically 12 seconds in our experiments). For the test data, as noted, experiments were conducted within the same domain as the initial dataset to later assess the effect of catastrophic forgetting upon domain expansion. Table III presents the training time and estimation results on both the initial and test datasets, while Figure 5 illustrates the model’s estimation results on the test data within domain $D_{1}$ following initial training.
@@
As shown in Table III and Figure 5, the model has an acceptable performance in terms of estimation error, as the maneuvers were performed in the same data domain as the initial dataset, there is no need to adapt the model to the test data.
@@ VI-B Phase 2: Continual Learning
Batch: Retrains all neural network parameters using both previously collected and newly acquired data, serving as a baseline or upper bound. This approach, referred to as Batch or Joint training mode.
@@
We consider a specified number of points as memory points for A-GEM ($m_{1}=150$), using the same quantity for EM-ReSeleCT. However, the key difference in EM-ReSeleCT is the selection of a specific representative set rather than random selection. This multivariate analysis within EM-ReSeleCT’s structure significantly reduces the number of memory points required. In our approach, after processing historical data with EM-ReSeleCT, the total number of memory points is reduced to $m_{1}=97$.
Table IV presents the training time and Mean Absolute Error (MAE) of each continual learning strategy on the new dataset after introducing the domain into the model. As shown, all strategies demonstrate strong performance in estimation error for the new maneuver. In terms of training time, the Batch mode requires substantial time due to its use of all historical data combined with the new data during training. Interestingly, despite the additional computations for memory selection in EM-ReSeleCT and A-GEM, they exhibit comparable training times. This efficiency is attributed to the significantly reduced final memory set in EM-ReSeleCT compared to A-GEM or random selection strategies. SI and None modes have the same training time, as neither retains memory points in the continual learning phase.
@@
A critical aspect of continual learning techniques is their capacity to preserve previously learned knowledge across domains encountered so far. Table V and Figures 7 and 8 display the estimation results on the test data (in $D_{1}$) using different continual learning strategies, following the introduction of a new domain (in $D_{2}$) to the model. As shown, the proposed method (EM-ReSeleCT) achieves an estimation error comparable to that of the Batch mode, which serves as our baseline. EM-ReSeleCT significantly outperforms other popular CL strategies, such as A-GEM and SI. This superior performance over A-GEM is attributed to EM-ReSeleCT’s capability to select informative memory points rather than relying on random selection, coupled with optimization modifications that enable positive backward transfer, enhancing learning on previously encountered data. SI, on the other hand, shows poor performance, as weight regularization methods have consistently demonstrated limited accuracy in these regression problems.
@@
Figure 7 presents the estimation results of EM-ReSeleCT in comparison to the Batch and None modes. As illustrated, for both longitudinal ($V_{x}$) and lateral ($V_{y}$) velocities, the proposed method demonstrates promising performance, closely aligning with the Batch mode but with significantly lower computational time. In specific regions, EM-ReSeleCT effectively preserves information, whereas in the None mode—where prior information is disregarded—the model performs poorly.
@@
As previously mentioned, EM-ReSeleCT enhances A-GEM’s optimization algorithm to better preserve historical data, as outlined in Eq. (9). Figure 9 illustrates the loss on memory points across training epochs in a continual learning scenario. Our algorithm not only prevents an increase in memory point loss but also reduces it over certain epochs, demonstrating improved retention of historical information. In addition to the modified optimization, EM-ReSeleCT selects key information by identifying informative data points rather than using A-GEM’s random selection, which further mitigates catastrophic forgetting across all historical data.
@@ VI-D Uncertainty Quantification with Conformal Prediction
$q_{1-\alpha}=\text{Quantile}_{1-\alpha}\{r_{1},r_{2},\dots,r_{n}\}.$ (20)
@@
$\text{PI}(x_{\text{test}})=\left[\widehat{f}(x_{\text{test}})-q_{1-\alpha},\widehat{f}(x_{\text{test}})+q_{1-\alpha}\right],$ (21)
where $\widehat{f}(x_{\text{test}})$ is the estimated output for input $x_{test}$, and $q_{1-\alpha}$ is the quantile calculated in Eq 20 This interval is guaranteed to cover the true target value $y_{\text{test}}$ with probability at least $1-\alpha$, under the assumption that the calibration set is exchangeable with the test data. In this study, we applied conformal prediction to estimate the prediction intervals for the outputs $V_{x}$ and $V_{y}$. Separate quantiles were computed for each output dimension, providing tailored uncertainty bounds for multivariate predictions.
To ensure reliable uncertainty quantification, we recalibrate the conformal prediction model after introducing the new dataset. The calibration data includes samples from both the old and new datasets to ensure comprehensive coverage. After determining the quantile $q_{\alpha}$ for each model, we apply the uncertainty quantification to both the test data from the old data space and the new data space.
Figure 10 (a) and (b), show the coverage rate and interval width for the estimation of $V_{x}$ with different numbers of memory points selected for the proposed model on the test dataset from the old data space.
Figure 10 (c) and (d), illustrate the coverage rate and interval width for the estimation of $V_{y}$ with different numbers of memory points selected for the proposed model.
As illustrated in the figures, the interval size and uncertainty level increase significantly as the number of memory points decreases. Furthermore, the figures demonstrate that after increasing the number of memory points beyond a certain threshold, the interval size remains unchanged. This suggests that storing additional memory points beyond this threshold does not contribute to further efficiency gains. The reason is that the memory points selected by the automatic approach discussed in the previous section act as representatives of historical data. Over-selecting leads to retaining redundant, non-informative samples that neither enhance information preservation nor effectively mitigate catastrophic forgetting.
Figure 10: Uncertainty analysis of different models by number of selected memory points, on the old (test) dataset. a and b: coverage rate and interval width vs the number of memory points on $Vx$, respectively. c and d: coverage rate and interval width vs the number of memory points on $Vy$, respectively.
@@ VII Conclusions and Future Work
In this paper, we introduce EM-ReSeleCT, an efficient multivariate representative selection and optimization strategy for continual learning in time series tasks. Experimental analysis on multivariate output estimation using Transformer models, applied to the electric Equinox vehicle, demonstrates the superior performance of EM-ReSeleCT in both training time and estimation accuracy compared to other state-of-the-art continual learning algorithms. Specifically, EM-ReSeleCT achieves estimation errors close to the Batch mode while significantly reducing computational time. The method effectively minimizes the number of memory points by identifying instances where two selected points across outputs represent the same information but are in close proximity.